Reflection of a plane wave at the interface between two dielectrics.

1
Expert's answer
2022-01-03T14:33:51-0500

We know that

Incidence wave "E=E_0e^{j(k.r-wt)}"

"Reflected wave 
\\E_1=E_{01}e^{j(k_1.r-w_1t)}"

Refracted wave

"E_2=E_{02}e^{j{(k_2.r-w_2t)}}"



"k_1" = Reflected wave propagation constant

"k_2" =Refracted wave propagation constant

"k=" Incidence wave propagation constant

"k_{1x}=k_{2x}=k_x"

"k_{1r}=k_{2r}=k_r"

For reflection

"k_{1x}=k_{x}"

"k_1sin\theta_1=ksin\theta"

"k_1=k"

"sin\theta_1=sin\theta"

"\theta_1=\theta"

This is reflection law